the tangent to the curve 𝑦 = 𝑥² − 5𝑥 − 2 at the point (1,-6) intersects the normal to


the same curve at the point (2, -8) at point P . find the coordinates of P

1
Expert's answer
2022-03-17T13:45:58-0400

Solution


Given that


"y = {x^2} - 5x - 2"


The gradient of the tangent is


"{m_t} = \frac{{dy}}{{dx}} = 2x - 5"


The gradient of the tangent at the point (1, - 6) is


"{m_t} = 2\left( 1 \right) - 5 = - 3"


Hence equation of the tangent at the point (1, -6) is


"\begin{gathered}
 y - {y_1} = {m_t}\left( {x - {x_1}} \right) \\
 y - \left( { - 6} \right) = - 3\left( {x - 1} \right) \\
 y + 6 = - 3x + 3 \\ 
\end{gathered}"


"3x + y + 3 = 0" .... (1)


Now we know the gradient of the tangent is "{m_t} = \frac{{dy}}{{dx}} = 2x - 5"


Therefore, the gradient of the normal will be "{m_n} = - \frac{1}{{{m_t}}} = - \frac{1}{{2x - 5}}"


And gradient of the normal at the point (2, -8) is "{m_n} = - \frac{1}{{2\left( 2 \right) - 5}} = - \frac{1}{{ - 1}} = 1"


Therefore, the equation of the normal at the point (2, -8) is 


"\begin{gathered}
 y - {y_1} = {m_n}\left( {x - {x_1}} \right) \\
 y - \left( { - 8} \right) = 1\left( {x - 2} \right) \\
 y + 8 = x - 2 \\ 
\end{gathered}"


"x - y - 10 = 0" .... (2)


Now the solving (1) and (2), we get 


"\begin{gathered}
 3x + y + 3 + x - y - 10 = 0\\
 4x - 7 = 0 \\
 x = \frac{7}{4} \\ 
\end{gathered}"


Replacing in (2)


"\begin{gathered}
 \frac{7}{4} - y - 10 = 0 & \\
 y = - \frac{{33}}{4} \\ 
\end{gathered}"  


Hence the point of intersection of the tangent and the normal is

 

"\left( {x,y} \right):\left( {\tfrac{7}{4},\,\,\tfrac{{ - 33}}{4}} \right) = \left( {1.75,\, - 8.25} \right)"


The tangent, normal, and the coordinates of the point P, where the tangent and the normal intersect are shown in the plot below.